Pré-Publicações do Departamento de Matemática Universidade de Coimbra Preprint Number 05–01

ACTUAL TITLE

AUTHOR 1, AUTHOR 2 AND AUTHOR 3 Dedicated to ourselves

Abstract: An abstract should be written using the environment defined as in this example file.

The text describes briefly the use of the prepdmuc class. Keywords: this environment is optional. AMS Subject Classification (2000): this environment is optional.

1. Introduction

This file gives a brief description of the class prepdmuc.cls. This class is based on the standard article class of the L^{A}TEX distribution. You have available the standard commands and environments as is the standard article class, with the changes in the preamble that will be described.

2. Preamble

As in the example file, the code starts with the instruction

\documentclass[a4paper]{prepdmuc}

possibly with the usual optional arguments. There should be no \pagestyle command as well as no length commands. You should provide a title and a running title. The instruction for this purpose is

\title[Running title]{Actual title}

You are expected to give two arguments, even if they coincide. Notice that the running title is to be used in the headers of the odd pages so it should not be long. For each author you should identify his/hers name, an abbreviated version of the name for the headers of even pages and possibly include some acknowledgments. Also you should give the address. If they exist and you wish to make them available, you may give the email and the url addresses. For this purpose you should use the following instructions:

Received August 1, 2005.

Thanks of Author 2 to someone.

Author 3 wishes to thanks himself.

\autor[abbrev. name]{Full name}{Address of author}{Any thanks.}
\email{@somewhere.pt}
\url{http://www.somewhere.pt/$\sim$author1}

Note that the \autor command expects two arguments, even if they coincide. In case of several authors, you should use, in the order given, the instructions concerning author 1, then author 2, and so on.

Include also the date of submission to the editor, with the command

\date{Month dd, year}

Besides these instructions, include your command definitions.

3. The document

3.1. Generating the title. As expected, the body of the document should start with the command \maketitle, which generates the title and the identification of the authors.

3.2. The abstract. There is an environment to handle the layout of the abstract. It goes, as expected, like this

\begin{abstract}
Abstract text.
\end{abstract}

3.3. Text. You do not need to change anything else in your file. All the usual L^{A}TEX sectioning commands are available. If including graphics or figures please attach the files.

3.4. References. Please use the \thebibliography environment with numbered or named references.

References

[1] B. Banaschewski, The real numbers in pointfree topology, Textos de Matemática, Série B, Vol. 12, Universidade de Coimbra, 1997.

[2] M.P. Fourman and J.M.E. Hyland, Sheaf models for analysis, Applications of Sheaves, Springer LNM 753 (1979) 280-301.

[3] R. Giles and H. Kummer, A non-commutative generalization of topology, Indiana Univ. Math. J. 21 (1971) 91-102.

[4] P.T. Johnstone, Stone Spaces, Cambridge Studies in Advanced Mathematics, Vol. 3, Cambridge University Press, Cambridge, 1982.

[5] P.T. Johnstone, The point of pointless topology, Bull. Amer. Math. Soc. 8 (1983) 41-53.

[6] A. Joyal and M. Tierney, An extension of the Galois Theory of Grothendieck, Mem. Amer. Math. Soc. 309, 1984.

[7] S. Vickers, Topology via Logic, Cambridge Tracts in Theoretical Computer Science, Vol. 5, Cambridge University Press, Cambridge, 1985.

Author 1 Department of Mathematics, University 1 of Sampleville, 1 Euler Avenue, Sampleville, UTOPIA

E-mail address: author1@math1.unisamp.ut
URL: http://www.math1.unisamp.ut/∼author1

Author 2 Department of Mathematics, University 2 of Sampleville, 2 GaußAvenue, U-314159 Sampleville, UTOPIA

E-mail address: author2@anywhere.pt

Author 3 Address of author 3

E-mail address: author3@here.pt
URL: http://www.here.pt/∼author3